What's the FCFF of a company with total revenues of $843 million, operating profit margin of 32%, tax rate of 31% and reinvestment rate of 37%? Answer in millions, rounded to one decimal place.
Answer #1

FCFF = 843*32%

=269.32,

NOW WE NEED TO PAY THE TAXES FIRST

=269.32*(1-.31)

=269.32*.69

=186.1344

NOW WE NEED TO DEDUCT THE REINVESTMENT

=186.1344* (1-.37)

=186.1344*.63

=117.26

FCFF =$117.3 MILLIONS

====